OpenStudy (anonymous):

For every real number x (regardless of whether x is positive, negative, or zero), sqrt(x^2)=x. True or false?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

\[\sqrt{x^2}=|x|\]

OpenStudy (anonymous):

so for example \[\sqrt{(-3)^2}=3\]

myininaya (myininaya):

gj satellite you are so smart you should reward yourself and right your bike